# Mills’ constant is irrational

### v1

@@ Abstract.
Let $\lfloor x\rfloor$ denote the integer part of $x$. In 1947, Mills constructed a real number $\xi$ greater than $1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$. We define Mills’ constant as the smallest real number $\xi$ satisfying this property. In this article, we determine that Mills’ constant is irrational. Moreover, we also obtain partial results on the transcendency of the constant.
@@ 1. Introduction
Let $\mathbb{N}$ denote the set of all positive integers. Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$. A complex number $\alpha$ is transcendental if $P(\alpha)\neq 0$ for all polynomials $P\in\mathbb{Q}[x]\setminus\{0\}$.
Mills [12] constructed a constant $\xi>1$ such that
@@
Such $\xi$ is not uniquely determined. In this article, we define Mills’ constant [^1] There are two types of definitions of Mills’ constant. Finch [7] implicitly introduced Mills’ constant as a real number $\xi$ satisfying (1.1). Our definition can be seen in [3]. as the smallest $\xi>1$ satisfying (1.1). It is well-defined from [13, Remark 4.2], which means that Mills’ constant exists. Assuming the Riemann hypothesis, Caldwell and Cheng [3] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately $1.3063778838\cdots$. It is an unsolved problem to determine whether the constant is rational or irrational. The goal of this article is to solve it.
Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi$ greater than $1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$. By [11, Corollary 4] and [13, Remark 4.2], $\xi_{c}$ exists for every integer $c\geq 2$. A real algebraic integer greater than $1$ is a Pisot number if all of its conjugates over $\mathbb{Q}$ except for itself lie in the open unit disc. We obtain the following results.
@@
For every integer $c\geq 4$, the constant $\xi_{c}$ is transcendental.
@@
Either the constant $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
For more generalization, let $(c_{k})_{k\in\mathbb{N}}$ be a real sequence with $c_{k}\geq 1$ for every $k\in\mathbb{N}$. We define $C_{k}=c_{1}\cdots c_{k}$ for every $k\in\mathbb{N}$ and
$\mathcal{W}(c_{k})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
An element $\xi\in\mathcal{W}(c_{k})$ is called a prime-representing constant (for short PRC). Wright [15] firstly investigated the geometric structure of $\mathcal{W}(c_{k})$. Especially, by his result [15, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$. Therefore, there exist infinitely many transcendental numbers in $\mathcal{W}(c_{k})$, but we do not get any arithmetic properties of the smallest element in $\mathcal{W}(c_{k})$ from this result. Alkauskas and Dubickas [1, Theorem 1] firstly gave a result on the transcendency of specific PRCs. They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if a sequence $(c_{k})_{k\in\mathbb{N}}$ of real numbers satisfying
@@
They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental or not. The author and Takeda [13, Theorem 1.3] showed that the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental if a sequence $(c_{k})_{k\in\mathbb{N}}$ of integers satisfying
@@
They also obtained a result on the algebraic independency of PRCs in [13]. The condition (iii) is essentially needed in these works, but we accomplish replacing (iii) with the condition $c_{k+1}\geq 4$ for infinitely many positive integers $k$ as follows.
@@
Then the smallest real number in $\mathcal{W}(c_{k})$ exists. Let $\xi$ be the number, and the following properties are true:
@@
Theorem 1.3 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.3 in Section 4. At the last of this section, let us ask several remaining questions.
Question 1.4.
@@
If the answer to Question 1.4 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2. In Section 5, we will give assumptions leading to the transcendency of $\xi_{3}$. We will explain the main difficulties of Questions 1.4 and 1.5 in Remarks 3.6 and 4.4, respectively.
2. Strategy of proofs
In this section, we demonstrate a strategy of proofs. We first recall Mills’ construction, that is, we now construct $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
Let $\mathcal{P}$ be the set of all prime numbers, and let $\theta=21/40$. We apply the following theorem given by Baker, Harman, and Pintz [2] [^2] Mills applied the result given by Ingham [8] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$. Thus, the construction of this section is modernly modified. .
@@ 2. Strategy of proofs
There exists a constant $d_{0}>0$ such that
@@
Since $23\theta=63/40<$ 2 and $p_{1}$ is sufficiently large, we see that
@@
and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}\eqqcolon w$ and $\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w^{\prime}$ exist [^3] Actually, $w=w^{\prime}$ is true, but we do not use this formula.. Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}\leq w^{\prime 3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
We will further give a proof of the following result for sharing the main idea.
@@
We actually want to discuss the arithmetic properties of $\xi_{3}(=\min\mathcal{W}(3))$. There is a gap between $w$ and $\xi_{3}$. However, this gap is not essential to prove the irrationality (and transcendency). We can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.1 and 3.2).
To prove Theorem 2.2, we estimate the least distance between $w^{3^{k}}$ and prime numbers.
@@
for every sufficiently large $k\in\mathbb{N}$. Therefore,
$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}.$
By applying the Maclaurin expansion, $(1+x)^{1/3}=1+O(x)$ holds if $|x|\leq 1$. Thus, since $\theta-1=-19/40<0$, for every sufficiently large $k$, we have
$p_{k}\leq w^{3^{k}}\leq p_{k}\left(1+O(p_{k}^{3(\theta-1)})\right)=p_{k}+O(p_{k}^{3(\theta-1)+1}),$
and hence there exists $K_{0}>0$ such that for every sufficiently large $k$
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}.$
Combining $03\theta-2=63/40-2<$ 0 and the most left-hand side of (2.1), we obtain
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}\leq K_{0}p_{k-1}^{3(3\theta-2)}\leq\cdots\leq K_{0}p_{1}^{3^{k-1}(3\theta-2)},$
which completes the proof of Lemma 2.4. ∎
Lemma 2.4 means that $w^{3^{k}}$ is very close to some prime number. On the other hand, the powers of a rational number are at a distance from integers. Indeed, in 1957, Mahler [9] proved the following result. Here, for every $x\in\mathbb{R}$, we define
@@
Let $r/s$ be a rational number greater than $1$ and which is not an integer. Let $\epsilon$ be a positive real number. Then, there exists an integer $n_{0}=n_{0}(r/s,\epsilon)$ such that
$\|(r/s)^{n}\|>e^{-\epsilon n}$
@@
We first assume that $w$ is an integer. Then
$p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3},$
which is a contradiction since $p_{2}$ is a prime number. Therefore, $w$ is not an integer. We secondly assume that $w$ is a rational number. Since $w$ is not an integer, there exist co-prime integers $r$ and $s$ with $r>s\geq 2$ such that $w=r/s$. By Lemma 2.4, there exists $\gamma>0$ such that for every sufficiently large $k$
$\|w^{3^{k}}\|\leq|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
On the other hand, by Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$, for every sufficiently large $k$
$\|w^{3^{k}}\|>e^{-3^{k}\gamma}.$
This is a contradiction. Therefore, $w$ is irrational. ∎
Through the discussion above, we can share the main idea of the proof of Theorem 1.3. We will compare the following two phenomena:
  • (i)
    powers of the smallest PRC are very close to some prime number;
  • (ii)
@@
For (ii), we need a generalization of Theorem 2.5 to real algebraic numbers. Corvaja and Zannier [4] gave this generalization by applying the $p$-adic Schmidt subspace theorem. Dubickas used their result and showed the following [6, Lemma 6] to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2).
@@
Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients satisfying its leading coefficient is positive. Wagner and Ziegler [14] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
@@
They showed that $\lim_{n\to\infty}x_{n}^{1/d^{n}}=\alpha>1$ exists and it is either irrational or an integer. Further, Dubickas [6] improved their result to the transcendency of $\alpha$ by applying Theorem 2.6. Their methods are very useful not only for the recursive equations (2.2) but also for recursive inequalities like (2.1). In addition, Dubickas presented the following lemma [6, Lemma 8].
@@
Let $\beta$ be a Pisot number of degree $\ell\geq 2$ with conjugates $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ labelled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. Then, there is a constant $\lambda>0$ depending only on $\beta$ such that
@@
3. Lemmas
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. By [13, Theorem 1.2], the set $\mathcal{W}(c_{k})$ is non-empty, closed, and lower bounded. Therefore, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists. We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by [13, Theorem 1.3]. Thus setting $B=\sup_{k\in\mathbb{N}}c_{k}$, we have $B<\infty$. Let $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$ which is an infinite set from (3). Let $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for all $k\in\mathbb{N}$. Let $\theta=21/40$. The following Lemmas 3.1 and 3.2 can be seen in [13, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.

### v2

@@ Abstract.
Let $\lfloor x\rfloor$ denote the integer part of $x$. In 1947, Mills constructed a real number $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$. We define Mills’ constant as the smallest real number $\xi$ satisfying this property. Determining whether this number is irrational has been a long-standing problem. In this paper, we show that Mills’ constant is irrational. Furthermore, we obtain partial results on the transcendency of this number.
@@ 1. Introduction
Let $\mathbb{N}$ denote the set of all positive integers. Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$. Mills [16] constructed a real number $\xi>1$ such that
@@
Such $\xi$ is not uniquely determined. In this paper, we define Mills’ constant [10, p.130] as the smallest $\xi>1$ satisfying (1.1). In Lemma 3.1, we will show the existence of the smallest element of such $\xi$’s. Assuming the Riemann hypothesis, Caldwell and Cheng [4] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately equal to $1.3063778838\cdots$. Proving the irrationality of a specific number satisfying (1.1) has remained an unsolved problem for over 70 years since Mills’ result. In this paper, we determine that Mills’ constant is irrational.
Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi>1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$. The number $\xi_{c}$ exists for every integer $c\geq 2$ by Corollary 3.4 (or [15, Corollary 4] and [18, Remark 4.2]). A real algebraic integer greater than $1$ is a Pisot number if all of its conjugates over $\mathbb{Q}$ except for itself lie in the open unit disc. We obtain the following results.
@@
For every integer $c\geq 4$, $\xi_{c}$ is transcendental.
@@
Either $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
For more generalization, let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers with $c_{k}\geq 1$ for all $k\in\mathbb{N}$, and we define
$\mathcal{W}(c_{k})=\mathcal{W}((c_{k})_{k=1}^{\infty})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\},$
where $C_{k}=c_{1}\cdots c_{k}$ for all $k\in\mathbb{N}$.
Remark 1.3.
We always use the notation $\mathcal{W}(c_{k})$ as $\mathcal{W}((c_{k})_{k=1}^{\infty})$ for simplicity, and so we remark that $\mathcal{W}(c_{k})$ is independent of $k$. For instance, if $(c_{k})_{k=1}^{\infty}=(3,3,3,\ldots)$, then
$\mathcal{W}(3)=\{\xi>1\colon\lfloor\xi^{3^{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
An element $\xi\in\mathcal{W}(c_{k})$ is called a prime-representing constant (for short PRC). After Mills’ result [16], there are several studies for extending this result. For example, in 1950, Kuıpers [12] constructed an element in $\mathcal{W}(c)$ for all integers $c\geq 3$. In 1951, Ansari [2] extended the range of $c$ to all real numbers $c>77/29=2.655\cdots$. In the same period, Niven [17] independently gave a similar extension, but he extended $c$ to real numbers $c>8/3=2.666\cdots$. Furthermore, in 1954, Wright [20] first studied the set of PRCs [^1] In honor of Wright, the author uses the symbol “$\mathcal{W}$” as the set of prime-representing constants. Wright actually considered a wider class of representing functions.. We refer the readers to Dudley’s survey [9] for more details. It is a good survey of the early research of Mills-type constants.
Especially, by Wright’s result [20, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$. Therefore, there exist infinitely many transcendental numbers in $\mathcal{W}(c_{k})$, but we do not get any arithmetic properties of the smallest element in $\mathcal{W}(c_{k})$ from this result. Alkauskas and Dubickas [1, Theorem 1] gave a result on the transcendency of specific PRCs. They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if the sequence $(c_{k})_{k=1}^{\infty}$ satisfies
@@
They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental. The author and Takeda [18, Theorem 1.3] showed the following result.
Theorem 1.4.
Let $(c_{k})_{k=1}^{\infty}$ be a sequence of positive integers satisfying
  1. (i)
    $c_{1}\geq 1$;
  2. (ii)
    $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$;
  3. (iii)
@@
Then, the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental.
The author and Takeda also obtained a result on the algebraic independency of PRCs in [18]. The condition (iii) is essentially needed in these works, but we accomplish replacing (iii) with the condition $c_{k+1}\geq 4$ for infinitely many positive integers $k$ as follows.
@@
Then the smallest real number in $\mathcal{W}(c_{k})$ exists, say $\xi$. For this $\xi$ the following is true:
@@
Theorem 1.5 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.5 in Section 4. At the end of this section, let us ask several remaining questions.
Question 1.6.
@@
If the answer to Question 1.7 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2. In Section 5, we will give assumptions leading to the transcendency of $\xi_{3}$. We will explain the main difficulty of Question 1.7 in Remark 4.4.
2. Strategy of our proof
In this section, we demonstrate a strategy of our proof. We first recall Mills’ construction, that is, we now construct $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
Let $\mathcal{P}$ be the set of all prime numbers, and let $\theta=21/40$. We apply the following theorem given by Baker, Harman, and Pintz [3, p.562] [^2] Mills applied the result given by Ingham [11] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$. Thus, the construction of this section is modernly modified. .
@@ 2. Strategy of our proof
There exists a real number $d_{0}>0$ such that
@@
By $23\theta=63/40<$ 2, we see that
@@
and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}=\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w$ exists. Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
We will further show the following result for giving a sketch of the proof.
@@
We actually want to discuss the arithmetic properties of $\xi_{3}(=\min\mathcal{W}(3))$. There is a gap between $w$ and $\xi_{3}$, but we can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.5 and 3.6).
To prove Theorem 2.2, we estimate the distance between $w^{3^{k}}$ and the nearest integer.
@@
for every $k\in\mathbb{N}$. Therefore,
$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}<p_{k}+2p_{k}^{3\theta-2},$
where the last inequality follows from $(1+x)^{1/3}<1+x$ for all $x>0$. Thus, for every $k\in\mathbb{N}$, we have
$|w^{3^{k}}-p_{k}|\leq 2p_{k}^{3\theta-2}.$
Combining $03\theta-2=63/40-2<$ 0 and the first inequality of (2.1), we obtain
$|w^{3^{k}}-p_{k}|\leq 2p_{k}^{3\theta-2}\leq 2p_{k-1}^{3(3\theta-2)}\leq\cdots\leq 2p_{1}^{3^{k-1}(3\theta-2)},$
which completes the proof of Lemma 2.4 by selecting $\gamma=\frac{2-3\theta}{6}\log p_{1}>0$. ∎
Lemma 2.4 means that $w^{3^{k}}$ is very close to some integer (especially, prime number). On the other hand, the powers of a rational number are at a distance from integers. Indeed, in 1957, Mahler [13] proved the following result. Here, for every $x\in\mathbb{R}$, we define
@@
Let $\alpha$ be a rational number strictly greater than $1$ and which is not an integer. Let $\epsilon$ be a positive real number. Then, there exists an integer $n_{0}=n_{0}(\alpha,\epsilon)$ such that
$\|\alpha^{n}\|>e^{-\epsilon n}$
@@
It is clear that $w\notin\mathbb{Z}$ since if not, then we have $p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3}$, a contradiction. Therefore, $w$ is irrational by combining Lemma 2.4 and Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$. ∎
Through the discussion above, we can give a sketch of the proof of Theorem 1.5. We will compare the following two phenomena:
  • powers of the smallest PRC are very close to some integer;
@@
For observing the second phenomenon, we need a generalization of Theorem 2.5 to real algebraic numbers. Corvaja and Zannier [6] gave this generalization by applying the $p$-adic Schmidt subspace theorem. Dubickas used their result and prepared the following theorem to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2) (see [8, Lemma 6]).
@@
Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients such that its leading coefficient is positive. Wagner and Ziegler [19] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
@@
They showed that $\lim_{n\to\infty}x_{n}^{1/d^{n}}=\alpha>1$ exists and it is either irrational or an integer. Further, Dubickas [8] improved their result to the transcendency of $\alpha$ by applying Theorem 2.6. Their methods are instrumental not only for the recursive equations (2.2) but also for recursive inequalities like (2.1). In addition, Dubickas presented the following lemma [8, Lemma 8].
@@
Let $\beta$ be a Pisot number of degree $\ell\geq 2$ with conjugates $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ labelled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. Then, there is a real number $\lambda>0$ depending only on $\beta$ such that
@@
3. Lemmas and auxiliary results
Lemma 3.1.
Let $(c_{k})_{k=1}^{\infty}$ be a sequence of positive real numbers with $c_{k}\geq 1$ for all $k\in\mathbb{N}$. If $\mathcal{W}(c_{k})$ is non-empty, then the smallest element in $\mathcal{W}(c_{k})$ exists.
Proof.
Since $\mathcal{W}(c_{k})$ is lower bounded and non-empty, $\inf\mathcal{W}(c_{k})$ exists. Let $\xi=\inf\mathcal{W}(c_{k})$. By the definition of the infimum, there is a sequence $A_{1}\geq A_{2}\geq\cdots$ such that $A_{j}\in\mathcal{W}(c_{k})$ for every $j\in\mathbb{N}$ and $\lim_{j\to\infty}A_{j}=\xi$. By the right-side continuity of the floor function, for all fixed $k\in\mathbb{N}$, we have
$\lfloor\xi^{C_{k}}\rfloor=\lfloor\lim_{j\to\infty}A_{j}^{C_{k}}\rfloor=\lim_{j\to\infty}\lfloor A_{j}^{C_{k}}\rfloor.$
Thus, we have $\lfloor\xi^{C_{k}}\rfloor\in\mathcal{P}$ since $\lfloor A_{j}^{C_{k}}\rfloor\in\mathcal{P}$ for all $j\in\mathbb{N}$ and $\lim_{j\to\infty}\lfloor A_{j}^{C_{k}}\rfloor$ converges in $\mathbb{Z}$. Therefore, $\inf\mathcal{W}(c_{k})=\xi\in\mathcal{W}(c_{k})$, which means that $\xi$ is the smallest element of $\mathcal{W}(c_{k})$. ∎
Theorem 3.2.
Let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers satisfying $c_{1}>0$ and $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$. Then, $\mathcal{W}(c_{k})$ is non-empty.
Proof.
See [15, Theorem 3] or [18, Theorem 3.3]. ∎
Remark 3.3.
In [15], Matomäki actually showed that $\mathcal{W}(c_{k})$ is uncountable, nowhere dense, and has Lebesgue measure 00 if $c_{k}\geq 2$ for all $k\in\mathbb{N}$. Thus, there is a small gap between the initial conditions $c_{1}>0$ in Theorem 3.2 and $c_{1}\geq 2$ in Matomäki’s result, but it is not essential by considering the locally bi-Lipschitz continuous map
$f_{\alpha}\colon\mathcal{W}(c_{1},c_{2},c_{3},\ldots)\ni A\longmapsto A^{1/\alpha}\in\mathcal{W}(\alpha c_{1},c_{2},c_{3},\ldots)$
for some fixed real number $\alpha>0$. We refer the readers to [18, Remark 3.4] for more details.
Corollary 3.4.
Let $(c_{k})_{k=1}^{\infty}$ be a sequence of real numbers satisfying $c_{1}>0$ and $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$. Then, the smallest element in $\mathcal{W}(c_{k})$ exists.
Proof.
It is clear from Lemma 3.1 and Theorem 3.2. ∎
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k=1}^{\infty}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.5. By Corollary 3.4, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists.
We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by Theorem 1.4. Thus, we set
(3.1) $B=\sup_{k\in\mathbb{N}}c_{k}<\infty,$
Furthermore, we let
(3.2) $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$
which is an infinite set from (3). Let $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for all $k\in\mathbb{N}$. Let $\theta=21/40$. The following Lemmas 3.5 and 3.6 can be seen in [18, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.